What is the pH of a solition that is 1M in each of monobasic and dibasic sodium phosphate?
Answer #1

monobasic sodium phosphate NaH2PO4 is the weak acid.

dibasic sodium phosphate Na2HPO4 is the conjugate base.

Ka = 6.2 x 10-8

pKa = -log(Ka)

pKa = -log(6.2 x 10-8)

pKa = 7.2

According to Henderson - Hasselbalch equation,

pH = pKa + log([conjugate base] / [weak acid])

pH = pKa + log([Na2HPO4] / [NaH2PO4])

pH = 7.2 + log(1 M / 1 M)

pH = 7.2 + log(1)

pH = 7.2 + 0

pH = 7.2

pH of solution is 7.2